OpenStudy (anonymous):

integrate the following: indefinite integral [x squareroot(1+2x)]dx

OpenStudy (anonymous):

u sub. let u = 1+2x

OpenStudy (anonymous):

once you have that, rewrite sqrt u = u^(1/2)

OpenStudy (anonymous):

should be able to integrate using only the pwoer rule

OpenStudy (anonymous):

do i need to find what x is in terms of u?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

once i do that then what? I'm confused on that part.

OpenStudy (anonymous):

so far i have (u-1)/2 * u^1/2 * dx

OpenStudy (anonymous):

u = 2x+1 => du = 2dx => x = (u-1)/2 so the integral becomes \[\frac{ 1 }{ 2 } \int\limits \left( \frac{ u-1 }{ 2 } \right)\sqrt{u}\,\,du=\frac{ 1 }{ 2 } \int\limits \left( \frac{ u-1 }{ 2 } \right)u^{\frac{ 1 }{ 2 }}\,\,du=\frac{ 1 }{ 4 } \int\limits \left( u^{\frac{ 3 }{ 2 }}-u^{\frac{ 1 }{ 2 }} \right)\,\,du\]

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

you're welcome!